OpenStudy (anonymous):

i need medals will help anyone!!!!! just comment on this post what u need help with

OpenStudy (anonymous):

suppose f(x)=2x-5 and g(x)= l3x-1l find the value of f(1)+g(2) how do i set this up? please help

OpenStudy (anonymous):

ok i do not know but ask @jabez177 or @dan815

OpenStudy (anonymous):

anyone else

OpenStudy (anonymous):

I will give u medal but I don't know how to

OpenStudy (anonymous):

or will give guidance on how to get answer

OpenStudy (anonymous):

What do u mean?

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

if i cant find it out i will find someone who can

OpenStudy (anonymous):

how many solutions does the equation have 2(2x+5)=4(x+3)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

@dan815 and @jabez177 will help u

OpenStudy (athecool):

Give me a medal and i'll give you one!!

OpenStudy (anonymous):

why wouldnt u just keep your medal

OpenStudy (anonymous):

anyone else have a question

OpenStudy (athecool):

um you mean so if i give other people medals i lose mine?

OpenStudy (athecool):

Hey i'll rate you a qualified helper if you rate me. How's that sounds?

OpenStudy (anonymous):

me

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

here

jabez177 (jabez177):

Hello @Jesuslover14, welcome to OpenStudy. The way that system works is you go through the questions bar and look for a question and certain subject you can help with. Then you will get medals. :)